in a party if 1 out of every 3 guests brought bouquets,1 out of 5 brought gift items,1 out of 6 gave cash gifts and the remaining 27 guest gave gift vouchers and none of them gave two things. how many people attended the party
Found 2 solutions by jorel555, brohi:
Answer by jorel555(1290)

Let n be the total amount of guests. Then:
n/3+n/5+n/6+27=n
10n+6n+5n+810=30n
9n=810
n=90
There were 90 total guests!!!!!!!!!!!!!

Answer by brohi(1)

Let n be the total amount of guests. Then:
n/3+n/5+n/6+27=n
10n+6n+5n+810=30n
9n=810
n=90
There were 90 total guests